OpenStudy (anonymous):

i'm asked to draw graph for following equality: |z-1|+|z+1|=3 note that: z=x+yi

OpenStudy (phi):

you can think of z - 1 as z - (1+0i) and | z - (1+0i) | is the distance between point z and 1+0i on the complex plane similarly | z +1| is the distance between z and -1 so this question is asking for the locus of points whose distance to 1 and -1 sum to a constant 3 maybe an ellipse?

OpenStudy (anonymous):

yes it is an ellipse, but i don't know how to mathematically prove it. i have instructions in this way: \[|z+1|=|(x+1)+yi|=\sqrt{(x+1)^{2}+y^2}\] \[|z-1|=|(x-1)+yi|=\sqrt{(x-1)^{2}+y^2}\] Now I somehow have to play around with this: \[\sqrt{(x+1)^{2}+y^2}+\sqrt{(x-1)^{2}+y^2}=3\] to get ellipse equation which is : \[\frac{ 4x^2 }{9 }+\frac{ 4y^2 }{ 5 } = 1\]

OpenStudy (phi):

by definition it is an ellipse http://www.mathsisfun.com/definitions/ellipse.html

OpenStudy (phi):

I assume you made progress ? \[ \sqrt{(x+1)^{2}+y^2}+\sqrt{(x-1)^{2}+y^2}=3 \\ \sqrt{(x+1)^{2}+y^2}= 3 - \sqrt{(x-1)^{2}+y^2} \] square both sides \[ (x+1)^{2}+y^2 = 9 + (x-1)^{2}+y^2 -6 \sqrt{(x-1)^{2}+y^2}\\ 6 \sqrt{(x-1)^{2}+y^2}= 9 + (x-1)^{2}+\cancel{y^2} - (x+1)^{2}-\cancel{y^2} \\ 6 \sqrt{(x-1)^{2}+y^2}= 9 + x^2 -2x + 1 - ( x^2 +2x +1)\\ 6 \sqrt{(x-1)^{2}+y^2}= 9 -4x \] square both sides \[ 36\left((x-1)^2 + y^2\right) = 81 +16x^2 -72x \\ 36\left(x^2 -2x +1 + y^2\right) = 81 +16x^2 -72x\\ 36 x^2 -72x + 36 +36y^2 =81 +16x^2 -72x\\ 20 x^2 + 36y^2= 45 \] which simplifies to \[ \frac{4}{9}x^2 + \frac{4}{5} y^2 =1 \]